Order: acetaminophen (Tylenol) elixir 10 mg/kg p.o. a. Child weight is 10 pounds. b. Available: acetaminophen (Tylenol) elixir 15 mg/ 5mL. c. How many mL of acetaminophen (Tylenol) will the nurse administer? 2. Order: amoxicillin (Amoxil) 100 mg p.o. every 6 hours. a. Child weighs 15 lbs. b. Safe range is 25 mg/kg/24 hours to 30 mg/kg/24 hours. c. What is the safe 24 hour range in mg? (Record using one decimal place). d. Is the ordered dose safe for a 24 hour period? 3. Order: digoxin (Lanoxin) 18 mcg IV bid. a. Child weighs 7 lbs. b. The safe range is 10mcg/kg/24 hours - 12 mcg/kg/24 hours. c. What is the safe 24 hour range in mcg? (Record using one decimal place). d. Is the ordered dose safe for a 24 hour period? e. If the medication is supplied 50 mcg/mL, how many mL will the nurse administer per dose? 4. Order: phenytoin (Dilantin) 40 mg p.o. every 8 hours. a. Child weighs 27 lbs. b. The safe range is 8 mg/kg/24 hours 10 mg/kg/24 hours c. What is the safe dose range in mg? (Record using one decimal place). d. Is the prescribed dose safe? e. If the medication is supplied 125 mg/5 mL, how many mL will the nurse administer per dose? (Record using one decimal place)
Answer #1

ANSWER.

1) Calculation of volume(ml) of acetaminophen required.

Formula

Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q).

Here

Ordered dose = 10 mg/kg.

Weight of the child = 10 lb.

Convert lb to kg for calculation

1 lb = 0.454 kg.

Weight of this patient = 10 * 0.454

= 4.5 kg.

Dose required for this patient = 10 * 4.5

= 45 mg.

Dose in hand(H) = 15 mg.

Quantity(Q) = 5 ml.

As per above formula

Volume(ml) of drug required = D/H * Q

= 45/15 * 5

= 3 * 5

= 15 ml.

2) c) Safe 24hr range in mg.

Ordered dose = 100 mg every 6h.

Safe dose range = 25 mg/kg/24hr - 30 mg/kg/24hr.

Weight of the patient = 15 lbs.

Convert lbs to kg for calculation

1 lbs = 0.454 kg.

Weight of the patient = 15 * 0.454

= 6.8 kg.

Dose(mg) range for this patient = 25 * 6.8 to 30 * 6.8

= 170 mg to 204 mg per 24 hr.

d) Ordered dose is safe or not.

Ordered dose = 100 mg every 6 hr.

Every 6 hr is 4 times a day.

So

Ordered dose = 100 * 4

= 400 mg/day or 24 hr.

Safe dose range for this patient = 170 mg to 204 mg per 24 hr.

* The ordered dose(400mg/24hr) is SAFE for this patient.

3) c) Calculation of safe dose range in mcg.

Ordered dose = 18 mcg bid.

Safe dose range = 10 mcg/kg/day - 12 mcg/kg/day.

Weight of the patient = 7 lbs

Convert lbs to kg for calculation

1 lbs = 0.454 kg.

So

Weight of the patient = 7 * 0.454

= 3.17 kg.

Safe dose range for this patient 24 hrs = 10 * 3.17 to 12 * 3.17

= 31.7 mcg to 38 mcg per 24 hr.

d) Safe dose range or not.

Ordered dose = 18 mcg bid.

bid is two times a day.

So

Ordered dose = 18 * 2

= 36 mcg/day.

Safe dose range for this patient = 31.7 mcg to 38 mcg per day.

* Ordered dose (36 mcg/day) is SAFE for this patient.

e) Volume(ml) of drug required per dose.

Formula

Volume(ml) = Desired dose/Dose in hand * Quantity.

Here

Desired dose(D) = 18 mcg.

Dose in hand(H) = 50 mcg.

Quantity(Q) = 1 ml.

As per above formula

Volume(ml) of drug required = D/H * Q

= 18/50 * 1

= 0.36 * 1

= 0.36

= 0.4 ml per dose.

4) Calculation of safe dose range.

Ordered dose = 40 mg every 8 hrs.

Weight of the patient = 27 lbs.

Convert lbs to kg for calculation

1 lbs = 0.454 kg

Weight of the patient = 27 * 0.454

= 12.3 kg.

Safe dose range = 8 mg/kg/24 hr to 10 mg/kg/24 hr.

Safe dose range required for this patient = 8 * 12.26 to 10 * 12.3

= 98 mg to 123 mg per 24 hr.

d) Prescribed dose is safe or not.

Ordered dose = 40 mg every 8 hr.

Every 8 hr is 3 times a day.

Ordered dose = 40 * 3

= 120 mg per day or 24 hr.

Safe dose range = 98 mg to 123 mg per 24 hr.

So

* Ordered dose is SAFE for this patient.

e) Calculation of volume(ml) of drug required per dose.

Desired dose(D) = 40 mg.

Dose in hand(H) = 125 mg.

Quantity(Q) = 1 ml.

As per above formula

Volume(ml) of drug required = D/H * Q

= 40/125 * 1

= 0.3 * 1

= 0.3 ml.